Keith leaves his house for Sarah’s house riding his bicycle at 8 miles per hour. At the same time, Sarah leaves her house heading toward Keith’s house walking 3 miles per hour. The system of equations that represents the distance, d, each is from Sarah’s house in h hours is given below:
d=8.25-8h
d=3h
what does the solution of this system represent?

Answer by mananth(16946)

d=8.25-8h
d=3h
3h=8.25-8h
11h=8.25
h=8.25/11
h=0.75
d=3*0.75= 2.25 miles
d is the distance they will cross each other. after 3/4 hours